Suppose that industries within a country are perfectly competitive and exhibit constant returns to scale. The labor-share of income is 60% and the capital to labor ratio is K/L = 4. How large must the real rental rate of capital be if real wages are $2?
Answer #1

the production function

MPL=

MPK=

Cost is minimized MPL/MPK=w/r

So, K/L=()*(w/r)

4=(40/60)*(2/r)

r=$0.33